************************************************************************* C*-NEWS 44 01*11*1995/30*06*1996 *************************************************************************

C*-96/P.3601 GOODEARL K.R. C*-Algebras of Real Rank Zero whose K_{)}’s are not Riesz Groups Department of Mathematics - Univ. California Santa Barbara CA 93106 (USA) E-MAIL : goodearl@math.ucsb.edu

C*-96/P.3602 BRATTELI O., JORGENSEN P.E.T. Endomorphisms of B(H) , II: Finitely Correlated States on O_{n} Mathematics Institute - Univ. Oslo - PB 1053-Blindern N-0316 Oslo (Norway) E-MAIL : bratteli@math.uio.no

jorgen@math.uiowa.edu

C*-96/P.3603 THOMSEN K. Limits of certain subhomogeneous C*-Algebras - I Matematisk Institut - Ny munkegade - 8000 Aarhus (Denmark)

C*-96/P.3604 NIELSEN K.E. THOMSEN K. Limits of Circle Algebras Matematisk Institut - Ny munkegade - 8000 Aarhus (Denmark)

C*-96/P.3605 BLECHER D.P. Factorizations in Universal Operator Spaces and Algebras Department of Mathematics - Univ. Houston Houston TX 77204-3476 (USA) E-MAIL : dblecher@math.uh.edu

C*-96/P.3606 BLECHER D.P. A Completely Bounded Characterization of Operator Algebras Department of Mathematics - Univ. Houston Houston TX 77204-3476 (USA) E-MAIL : dblecher@math.uh.edu C*-96/P.3607 BLECHER D.P. A Generalization of Hilbert Modules (final revised version) Department of Mathematics - Univ. Houston Houston TX 77204-3476 (USA) E-MAIL : dblecher@math.uh.edu

C*-96/P.3608 BLECHER D.P. Continuous functions on Compact Groups Department of Mathematics - Univ. Houston Houston TX 77204-3476 (USA) E-MAIL : dblecher@math.uh.edu

C*-96/P.3609 BLECHER D.P. Le Merdy C. On Quotients of Function Algebras Department of Mathematics - Univ. Houston Houston TX 77204-3476 (USA) E-MAIL : dblecher@math.uh.edu C*-96/P.3610 EXEL R., LACA M. Continuous Fell Bundles Associated to Measurable Twisted Actions Departamento de Matematica - Univ. Sao Paulo - C.P. 20570 01452-990 Sao Paulo (Brazil) E-MAIL : exel@ime.usp.br

marcelo@math.newcastle.edu.au

C*-96/P.3611 GOTO S. Commutative of Automorphisms of Subfactors Modulo Inner Automorphisms Department of Mathematical Sciences - Univ. Tokyo - Hogo - Tokyo 113 (Japan) E-MAIL : s-goto@hoffman.cc.sophia.ac.jp

C*-96/P.3612 PHILLIPS N.C. A Classification Theorem for Nuclear Purely Infinite Simple C*-Algebras Department of Mathematics - Univ. Oregon Eugene OR 97403-1222 (USA)

C*-96/P.3613 PHILLIPS N.C.

Approximate Unitary Equivalence of Homomorphisms from Odd Cuntz Algebras Department of Mathematics - Univ. Oregon Eugene OR 97403-1222 (USA)

C*-96/P.3614 LIN H., PHILLIPS N.C. Almost Multiplicative Morphisms and the Cuntz Algebra O_{2} Department of Mathematics - Univ. Oregon Eugene OR 97403-1222 (USA)

C*-96/P.3615 MURPHY G.J., PHILLIPS N.C. C*-Algebras with the Approximate Positive Factorization Property Department of Mathematics - Univ. Oregon Eugene OR 97403-1222 (USA)

C*-96/P.3616 LIN H., PHILLIPS N.C. Classification of Direct Limits of Even Cuntz-Circle Algebras Department of Mathematics - Univ. Oregon Eugene OR 97403-1222 (USA)

C*-96/P.3617 LIN H., PHILLIPS N.C. Approximate Unitary Equivalence of Homomorphisms from O_{2} Department of Mathematics - Univ. Oregon Eugene OR 97403-1222 (USA)

C*-96/P.3618 RENAULT J. The Fourier Algebra of a Measured Groupoid and its Multipliers Département de Matématiques - Univ. orláns - B.P. 6759 45067 Orléans (France) E-MAIL : renault@univ-orleans.fr

C*-96/P.3619 JORGENSEN J.E.T. Harmonic Analysis of Fractal Processes via C*-Algebras Dpt. Math., Univ. Iowa - Iowa City IA 52242 (USA) E-MAIL : jorgen@math.uiowa.edu

C*-96/P.3620 SYMANSKI W., ZHANG S. Infinite Simple C*-Algebras and Reduced Cross Products of Abelian C*-Algebras and Free Groups Department of Mathematics, Univ. Newcastle, Newcastle NSW 2300 (Australia) E-MAIL : wojciech@frey.newcastler.edu.au

zhang@ucbeh.san.uc.edu

C*-96/P.3621 VAN DAELE A. An Algebraic Framework for Group Duality Department of Mathematics, Univ. Oslo (Norway)

C*-96/P.3622 VAN DAELE A. Multiplier Hopf Algebras and Duality Department of Mathematics, Univ. Oslo (Norway)

C*-96/P.3623 QUIGG J. Crossed Product Duality for Partial C*-Automorphisms Department of Mathematics, Arizona State Univ., Tempe AR 85287 (USA) E-MAIL : quigg@math.la.asu.edu

C*-96/P.3624 RADULESCU F. The Fundamental Group of the Von Neumann Algebra of any Cocompact Subgroup in PSL(2,R) Contains Q+/0 Harmonic Analysis of Fractal Processes via C*-Algebras Dpt. Math., Univ. Iowa - Iowa City IA 52242 (USA)

C*-96/P.3625 RADULESCU F. Analyse Fonctionnelle/Functional Analysis - Arithmetic Hecke Operators as Completely Positive Maps Harmonic Analysis of Fractal Processes via C*-Algebras Dpt. Math., Univ. Iowa - Iowa City IA 52242 (USA)

C*-96/P.3626 RADULESCU F. On the Rho-Equivariant Form of the Berezin’s Quantization of the Upper Half Plane Harmonic Analysis of Fractal Processes via C*-Algebras Dpt. Math., Univ. Iowa - Iowa City IA 52242 (USA)

C*-96/P.3627 DELAROCHE C.A. C*-Algebres purement infinies et groupes hyperboliques Département de Mathématiques - BP 6759 - Univ. Orléans 45067 Orléans Cedex 2 (France)

C*-96/P.3628 STORMER E. States and Shifts on Infinite Free Products of C*-Algebras Institute of Mathematics - Univ. Oslo - P.O. Box 1053 Blindern N-0316 OSLO (Norway)

C*-96/P.3629 FRANK M. Isomorphisms of Hilbert C*-Modules and *-Isomorphisms of Related Operator C*-Algebras Univ. Leipzig - Fak. Mat. Inf. Institut - Augustusplatz 10/11 04109 Leipzig (Germany)

C*-96/P.3630 DONSIG A.P., HOPENWASSER A. Order Preservation in Limit Algebras Department Mathematics - Univ. Lancaster - Lancaster LA1 4YF (U.K.) E-MAIL : a.donsig@lancaster.ac.uk

ahopenwa@ua1vm.ua.edu

C*-96/P.3631 EFFROS E.G., WEBSTER C. Operator Analogues of Locally Convex Spaces Mathematics Department - UCLA - Los Angeles CA 90024 (USA) E-MAIL : ege@math.ucla.edu

cwebster@math.ucla.edu

C*-96/P.3632 VALETTE A. On the Haagerup Inequality and Groups Acting on A_{n}-Building Institut de Mathématiques - Rue Emile Argand 11 CH-2007 Neuchâtel (Suisse) E-MAIL : valette@maths.unine.ch

C*-96/P.3633 KAWAHIGASHI Y. Classification of Approximately Inner Automorphisms of Subfactors Department Mathematical Sciences - Univ. Tokyo - Hongo Tokyo 113 (Japan)

E-MAIL : yasuyuki@tansei.cc.u-tokyo.ac.jp

C*-96/P.3634 SATO N. Fourier Transform for Paragroups and its Application to the Depth Two Case Department Mathematical Sciences - Univ. Tokyo - Hongo Tokyo 113 (Japan) E-MAIL : nn57012@hongo.ecc.u-tokyo.ac.jp

C*-96/P.3635 MCCANN P.J., CAREY A.L. A Discrete Model of the Integer Quantum Hall Effect Department of Pure Mathematics - Univ. Adelaide (Australie)

C*-96/P.3636 EFFROS E.G., WINKLER S. Matrix Convexity : Operator Analogues of the Bipolar and Hahn-Banach Theorems Mathematics Department - UCLA - Los Angeles CA 90024 (USA)

C*-96/P.3637 DADARLAT M., LORING T.A. A Universal Multi-Coefficient Theorem for the Kasparov Groups Department of Mathematics and Statistics - Purdue Univ. West Lafayette IN 47907 (USA)

C*-96/P.3638 HADWIN D., LORING T.A. Normal Operators in C*-Algebras without Nice Approximants Department of Mathematics and Statistics - Purdue Univ. West Lafayette IN 47907 (USA)

C*-96/P.3639 FRANK M., TROITSKY E. Leftchetz Numbers and Geometry of Operators in W*-Modules Univ. Leipzig ‘ Fak. Mat. Inf. Institut - Augustusplatz 10/11 04109 Leipzig (Germany)

C*-96/P.3640 KATAYAMA Y., SUTHERLAND C.E.,

TAKESAKI M.

The Characteristic Square of a Factor and the Cocycle Conjugacy of Discrete Group Actions on Factors Department of Mathematics - Univ. California - Los Angeles 90095-1555 (USA)

C*-96/P.3641 FRIIS P., RORDAM M. Almost Commuting Self-Adjoint Matrices - A Short Proof of Huaxin Lin’s Theorem Institut for Matematik of Datalogi - Odense Univ. Campusvej 55 - DK-5230 Odense (Denmark) E-MAIL : friis@math.toronto.edu

mikael@imada.ou.dk

C*-96/P.3642 GOTO S. Quantum Double Construction for Subfactors Arising from Periodic Commuting Squares Department Mathematics - Sophia Univ. 7-1 Kioi-cho - Chiyoda-ku - Tokyo 102 (Japan) E-MAIL : s-goto@hoffman.cc.sophia.ac.jp

C*-96/P.3643 SATO N. Two Subfactors Arising from a Non-Degenerate Commuting Square II. Tensor Categories and TQFT’s Dept. Mathematical Sciences - Univ. Tokyo - Komaba - Tokyo 153 (Japan) E-MAIL : nobuya@ms.u-tokyo.ac.jp

C*-96/P.3644 MASUDA T. An Analogue of Longo’s Canonical Endomorphism for Bimodule Theory and its Application to Asymptotic Inclusions Dept. Mathematical Sciences - Univ. Tokyo - Komaba - Tokyo 153 (Japan) E-MAIL : masuda@ms.u-tokyo.ac.jp